OpenStudy (anonymous):

find the derivative of g(x) = f(x)^3 by using the definition

OpenStudy (whpalmer4):

\[\lim_{h\rightarrow0}\frac {f(x+h)-f(x)}{h} =\lim_{h\rightarrow0}\frac {(x+h)^3-x^3}{h} = \lim_{h\rightarrow0}\frac {(x^3+3hx^2+3h^2x+h^3)-x^3}{h} \] You should be able to do the rest...

OpenStudy (anonymous):

@whpalmer4, what you posted was the derivative of \[f(x)=x^3,\] whereas the question was finding the derivative of \[g(x) = [f(x)]^3\] using the limit definition.

OpenStudy (anonymous):

According to the definition, the derivative of g(x) is \[\lim_{h\to 0} \dfrac{[f(x+h)]^3-[f(x)]^3}{h},\] which becomes much trickier to figure out.